A traffic study of 2,500 vehicles that passed by a checkpoint showed that their speeds were normally distributed with a mean of 82.5 kph and a standard deviation of 11.5 kph. How many vehicles had a speed of more than 90 kph


1
Expert's answer
2021-11-11T11:42:56-0500

"\bar x = 82.5"

"\sigma = 11.5"

"P(X>90)=1-P(X<90) \\
=1-P(Z<\frac{90-82.5}{11.5})"

"=1-P(Z<\frac{90-82.5}{11.5}) \\
=1-P(Z<0.652) \\
=1-0.7427\\
=0.2573"

25.73% of 2500 vehicles:

"2500 \times 0.2573= 643"

Answer: 643 vehicles have a speed of more than 90 kph.